OpenStudy (jazzyfa30):

Write the equation y = 2x +3 in standard form.

OpenStudy (uri):

y=2x+3 2x=y-3 2x-y=-3

OpenStudy (jazzyfa30):

can u help me with a few more

OpenStudy (uri):

Sure.

OpenStudy (uri):

No idea.

OpenStudy (jazzyfa30):

go can u help me

OpenStudy (goformit100):

Ya sure... but in one condition

OpenStudy (jazzyfa30):

what??

OpenStudy (goformit100):

Post the pics in More Bigger size

OpenStudy (jazzyfa30):

ok

OpenStudy (jazzyfa30):

@goformit100 come on